OpenStudy (anonymous):

The Frostburg-Truth bus travels on a straight road from Frostburg Mall to Sojourner Truth Park. The mall is 4 miles east and 4 miles north of the City Center. The park is 2 miles west and 4 miles south of the Center. How far is it from the mall to the park to the nearest tenth of a mile?

OpenStudy (anonymous):

First figure out how far North and East the mall is to the Park.

OpenStudy (nurali):

Then draw on the details you know. The Mall is 4 units East and 4 North of the City. The Park is 4 units South and 2 West of the City. So the Mall is, in total, 8 units North of the Park, and 6 units East of the park. There's your triangle. Sides 6 and 8 and it's a right-angled triangle. a² = b² + c² a² = 6² + 8² a² = 36 + 64 a = √100 a = 10...

OpenStudy (jdoe0001):

bubbles10021 read what Nurali said, and check the picture above :)